# Ranked from Within: Ranking Large Multimodal Models for Visual Question Answering Without Labels

### v1

@@ 7 Conclusion
This work studies an interesting question: ranking large multimodal models without the access to target domain labels. Our analysis begins by identifying weak correlations in model performance across different domains, leading us to explore uncertainty scores estimated from model predictions. We evaluate 45 LMMs on visual question answering, testing various training frameworks, visual encoders, and language models. Our experiments reveal that negative log likelihood serves as a useful performance indicator, and while stochastic sampling can be helpful, it is less effective for multiple-choice tasks, where it requires repeated inferences and careful temperature tuning. This study establishes a baseline for using uncertainty in model ranking, paving the way for further research
Several exciting directions remain open. Beyond uncertainty scores, test-time augmentation and semantic entropy may offer additional insights. Our findings suggest that text feature dissimilarity has a stronger impact on weak performance correlations across datasets than image feature dissimilarity. Exploring other factors to design more effective multimodal benchmarks presents an intriguing avenue.

### v2

@@ 7 Conclusion and Discussion
This work studied whether the performance of large multimodal models in a new target domain can be ranked without the access to target domain labels. Our analysis identified only a weak correlation in model performance across different domains. This motivated the investigation and evaluation of alternative approaches based on uncertainty estimates obtained from model predictions. We evaluate 45 LMMs on closed and open-ended visual question answering tasks, testing various training frameworks, visual encoders, and language models. Our experiments reveal that scores based on the negative log-likelihood of generated tokens serve as highly effective performance indicators for target domains. We also find that while stochastic sampling can be helpful, it is less effective for multiple-choice tasks, where it requires many repetitions of inference and careful temperature tuning. By establishing a baseline for uncertainty-based LMMs ranking, this study aims to motivate and inspire further research into this important area.
Potential future directions.
Beyond uncertainty scores, several promising directions remain for future research. Test-time augmentation and semantic entropy present natural next steps. Additionally, the internal states of LMMs (Azaria & Mitchell, 2023; Gottesman & Geva, 2024) offer opportunities for uncertainty estimation or error detection, for example, by training a classifier on these representations. However, in unsupervised LMM ranking, this approach requires training a separate classifier per model, which can be computationally expensive. An alternative is to measure the statistical distance between a model’s internal state for a given response and the distribution of internal states from multiple inference passes. A larger average distance may signal greater uncertainty and potentially indicate a lower model rank. Exploring how internal states can inform LMM ranking is an open direction. Finally, the observed asymmetry in the impact of text and image feature dissimilarity on cross-dataset correlations highlights a valuable area for improving multimodal benchmark design.
Acknowledgments
We thank all anonymous reviewers for their constructive feedback, which helped improve the quality of this paper. Tongliang Liu is partially supported by the following Australian Research Council projects: FT220100318, DP220102121, LP220100527, LP220200949, and IC190100031.
Impact Statement
This paper aims to advance machine learning while acknowledging potential societal impacts. Our findings could be misused: adversarial researchers may train models that consistently get ranked higher than other models, despite performing poorly on data. To mitigate this, incorporating robust calibration methods would be helpful, making model confidence accurately reflects uncertainty. This would help promote safe and responsible deployment of our findings.
Appendix A Appendix
In this supplementary material, we first introduce the experimental details including the models, datasets and compute in Appendix B. Next, we visualize the image and text features using t-SNE plots. Last, we show the full results of the correlation study on all datasets.
Appendix B Experiment Details
B.1 Datasets
All experiments are conducted on VLMEvalKit (Duan et al., 2024). We consider datasets and their corresponding links to download TSV files via the toolkit:
ScienceQA (Lu et al., 2022) (https://opencompass.openxlab.space/utils/VLMEval/ScienceQA_TEST.tsv);
AI2D (Hiippala et al., 2021) (https://opencompass.openxlab.space/utils/VLMEval/AI2D_TEST.tsv);
ChartQA (Masry et al., 2022) (https://opencompass.openxlab.space/utils/VLMEval/ChartQA_TEST.tsv);
OCRVQA (Mishra et al., 2019) (https://opencompass.openxlab.space/utils/VLMEval/OCRVQA_TESTCORE.tsv);
TextVQA (Singh et al., 2019) (https://opencompass.openxlab.space/utils/VLMEval/TextVQA_VAL.tsv);
DocVQA (Mathew et al., 2021) (https://opencompass.openxlab.space/utils/VLMEval/DocVQA_VAL.tsv);
RealWorldQA (x.ai, 2024) (https://opencompass.openxlab.space/utils/VLMEval/RealWorld.tsv);
MMMU (Yue et al., 2024) (https://opencompass.openxlab.space/utils/VLMEval/MMMU_DEV_VAL.tsv)
GQA (Ainslie et al., 2023) (https://opencompass.openxlab.space/utils/VLMEval/GQA_TestDev_Balanced.tsv)
B.2 Models
We include a diverse array of large multimodal models from series:
B.3 Compute and Library
PyTorch version is 2.01.0+cu117. All experiment is run on four A6000 GPUs. All models can be downloaded via Huggingface with different versions of transformer library:
Appendix C Visualization of Image and Text Features
In the main paper, we demonstrate that distances in text features contribute more significantly to the weak correlation of model performance across datasets than image features. To visualize this finding, we utilize t-distributed stochastic neighbor embedding (t-SNE) (Hinton & Roweis, 2002). Our results show that CLIP-L-14-336 (Radford et al., 2021) effectively captures distinctions between images from different datasets, with images from the same dataset forming distinct clusters. Additionally, the text features of visual question answering (VQA) and multiple-choice visual questioning (MCVQ) tasks are separated by the text encoder of CLIP. Notably, the text features of OCRVQA are distant from those of ChartQA, TextVQA, and DocVQA, despite all being VQA tasks. This finding supports our observation in the main paper that the effectiveness of ranking methods is influenced by dataset characteristics (i.e., VQA vs. MCVQ).
Appendix D Full Results of Correlation Study
In the following, we present the full results of correlation study for ranking different series of large multimodal models. We only show NLL, NLL, NLL, NLL, Ent, Ent, Ent, Ent, Sample, Sample, Sample*. ATC and Accuracy on the line are not included because their performance are computed by the average correlation strength across eight datasets.